# On the minimax optimality of Flow Matching through the connection to kernel density estimation

### v1

@@ 4.2 Proofs and extended results of Section 3
Using [Villani, 2008, Remark 6.5] we get
@@
Like in [Albergo and Vanden-Eijnden, 2022, Proposition 3] set
@@
Now we use the formulation of Grönwall’s lemma in Walter [1970] and use $Q_{0}=0$ (this holds since $\psi_{0}(x)=\hat{\psi}_{0}(x)=x$). This leads to
@@
By Theorem 2.2 the vector field $\hat{v}$ is chosen such that it minimizes $\int_{0}^{1}\mathbb{E}_{X_{t}\sim p_{t}}\big{[}|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big{]}\;\mathrm{d}t$ and $X_{t}=\psi_{t}(Z),$ thus we get for every $\tilde{v}\in\mathcal{M}$
@@
Using Youngs inequality and the properties of convolution, we obtain for $k\in\mathbb{N}^{d}_{0}$
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\leq\|D^{1}f\|_{\infty}\|D^{k-1}K_{\sigma_{\min}(-\cdot)}\|_{1}\leq 1\cdot\|D^{k-1}K_{\sigma_{\min}}\|_{1}.$
By assumption $\|D^{k-1}K\|_{1}\leq C$ and hence $\|D^{k-1}K_{\sigma_{\min}}\|_{1}\leq C\sigma_{\min}^{-k+1}.$ Now we can use [Schreuder, 2020, Theorem 4]. For even $d,$ set $k=\frac{d}{2}.$ Then there exists a constant $C^{\prime}_{2}>0$ such that
@@
$\frac{D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(x)-D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(y)}{|x-y|^{1/2}}\leq 2C^{\prime}_{2}\sigma^{-\frac{d-3}{2}}\leq 2C^{\prime}_{2}\sigma^{-\frac{d}{2}+1}.$
Then we can use [Schreuder, 2020, Theorem 4] with $\alpha=\frac{d}{2}$ again. Therefore, there is a $C^{\prime\prime}_{2}>0$ such that
$\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\leq\frac{C^{\prime\prime}_{2}}{\sigma^{\frac{d}{2}-1}}n^{-1/2}\log(n).$
@@
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sigma^{\frac{d}{2}-1}}n^{-1/2}\log(n).\qed$
Proof of Theorem 3.6
The proof of Theorem 3.6 requires some additional results. In order to control the error $e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\hat{v}\|_{\infty}$ (where we omitted constants) we need to approximate the function $v^{n}$ and control the Lipschitz constant of the apprximation $\hat{v}_{t}$ for fixed $t.$ However, we cannot expect to approximate a non-tivial function on $\mathbb{R}^{d}$ with respect to the supremum norm. Additionally, $v^{n}_{t}$ as constructed by Lipman et al. [2023] is only locally Lipschitz in $x$ for fixed $t$ and the bound of the Lipschitz constant grows for $\sigma_{\min}\rightarrow 0$ (see Lemma 4.2). Hence we cannot expect to construct a network $\hat{v}$ that approximates $v^{n}$ and its Lipschitz constant on $\mathbb{R}^{d}$ with a small absolute error for every $x\in\mathbb{R}.$

### v2

@@ 5.2 Proofs and extended results of Section˜3
Using (Villani, 2008, Remark 6.5) we get
@@
Like in (Albergo and Vanden-Eijnden, 2023, Proposition 3) set
@@
Now we use the formulation of Grönwall’s lemma in Walter (1970) and use $Q_{0}=0$ (this holds since $\psi_{0}(x)=\hat{\psi}_{0}(x)=x$). This leads to
@@
By Theorem˜2.2 the vector field $\hat{v}$ is chosen such that it minimizes $\int_{0}^{1}\mathbb{E}_{X_{t}\sim p_{t}}\big[|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big]\;\mathrm{d}t$ and $X_{t}=\psi_{t}(Z),$ thus we get for every $\tilde{v}\in\mathcal{N}$
@@
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $=\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
@@
For any multi index $k\in\mathbb{N}^{d}_{0}$ with $|k|\geq 1$, choose an index $j$ with $k_{j}\geq 1$ and set $k=\ell+e_{j}$ where $|k|=|\ell|+1$. Then using Young’s inequality and the properties of convolution, we obtain
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\leq\|D^{e_{j}}f\|_{\infty}\|D^{\ell}K_{\sigma_{\min}(-\cdot)}\|_{1}\leq 1\cdot\|D^{\ell}K_{\sigma_{\min}}\|_{1}.$
By assumption $\|D^{\ell}K\|_{1}\leq C$ and hence $\|D^{\ell}K_{\sigma_{\min}}\|_{1}\leq C\sigma_{\min}^{-|k|+1}.$ Thus
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\lesssim\sigma_{\min}^{-|k|+1}.$ (21)
Now we can use Schreuder (2020, Theorem 4). For even $d,$ set the smoothness parameter to $\frac{d}{2}.$ Then there exists a constant $C^{\prime}_{2}>0$ such that
@@
$\frac{D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(x)-D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(y)}{|x-y|^{1/2}}\leq 2C^{\prime}_{2}\sigma_{\min}^{-\frac{d-3}{2}}\leq 2C^{\prime}_{2}\sigma_{\min}^{-\frac{d}{2}+1}.$
Then we can use Schreuder (2020, Theorem 4) with $\alpha=\frac{d}{2}$ again. Therefore, there is a $C^{\prime\prime}_{2}>0$ such that
$\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\leq\frac{C^{\prime\prime}_{2}}{\sigma_{\min}^{\frac{d}{2}-1}}n^{-1/2}\log(n).$
@@
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sigma_{\min}^{\frac{d}{2}-1}}n^{-1/2}\log(n).\qed$
Proof of Theorem˜3.7
The proof of Theorem˜3.7 requires some additional results. The proof of the subsidiary lemmata can be found in Section˜5.3. In order to control the error $e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\hat{v}\|_{\infty}$ (where we omitted constants) we need to approximate the function $v^{n}$ and control the Lipschitz constant of the approximation $\hat{v}_{t}$ for fixed $t.$ However, we cannot expect to approximate a non-trivial function on $\mathbb{R}^{d}$ with respect to the supremum norm. Additionally, $v^{n}_{t}$ as constructed by Lipman et al. (2023) is only locally Lipschitz in $x$ for fixed $t$ and the bound of the Lipschitz constant grows for $\sigma_{\min}\rightarrow 0$ (see Lemma˜5.2). Hence we cannot expect to construct a network $\hat{v}$ that approximates $v^{n}$ and its Lipschitz constant on $\mathbb{R}^{d}$ with a small absolute error for every $x\in\mathbb{R}.$